# Prometheus: Infrastructure Security Posture Analysis with AI-generated Attack Graphs

### v1

@@ Abstract.
The rampant occurrence of cybersecurity breaches imposes substantial limitations on the progress of network infrastructures, leading to compromised data, financial losses, potential harm to individuals, and disruptions in essential services. The current security landscape demands the urgent development of a holistic security assessment solution that encompasses vulnerability analysis and investigates the potential exploitation of these vulnerabilities as attack paths. In this paper, we propose Prometheus, an advanced system designed to provide a detailed analysis of the security posture of computing infrastructures. Using user-provided information, such as device details and software versions, Prometheus performs a comprehensive security assessment. This assessment includes identifying associated vulnerabilities and constructing potential attack graphs that adversaries can exploit. Furthermore, Prometheus evaluates the exploitability of these attack paths and quantifies the overall security posture through a scoring mechanism. The system takes a holistic approach by analyzing security layers encompassing hardware, system, network, and cryptography. Furthermore, Prometheus delves into the interconnections between these layers, exploring how vulnerabilities in one layer can be leveraged to exploit vulnerabilities in others. In this paper, we present the end-to-end pipeline implemented in Prometheus, showcasing the systematic approach adopted for conducting this thorough security analysis.
@@ 1. Introduction
Our Approach. To address such challenges, we design an innovative solution, called Prometheus, which serves as a comprehensive security posture analyzer for computing infrastructures and applications. Prometheus operates seamlessly by continuously monitoring trustworthy data sources, such as national vulnerability databases, for disclosed vulnerabilities specific to the devices and configuration of a given infrastructure.
@@
Using named entity recognition (NER), Prometheus is able to extract the semantic meaning of these vulnerabilities and encode them into a latent space for in-depth analysis. In particular, Prometheus automatically extracts the preconditions required for an adversary to exploit a vulnerability and the result after exploiting the vulnerability (i.e., postconditions). For example, Prometheus can discern that exploiting a particular CVE might require the presence of the TensorFlow XLA compiler in the Google TensorFlow version prior to 1.7.0 as a precondition. Simultaneously, Prometheus captures postconditions such as a system crash resulting from a heap buffer overflow or reading from other parts of the process memory.
@@
Subsequently, Prometheus employs a semantic similarity-matching approach based on word embeddings (li2018word, ) to evaluate whether the postconditions of a vulnerability match the preconditions of another, thus constructing potential attack graphs for a given topology. Once generated, Prometheus analyzes these attack graphs to evaluate the security posture. We propose a novel layered approach to this analysis, where vulnerabilities within each layer share similar characteristics and can be comprehensively assessed by subject matter experts. Prometheus categorizes vulnerabilities in the generated attack graphs into distinct layers, such as machine learning, system, hardware, network, and cryptography, using keyword matching and common weakness enumeration (CWE) information within the vulnerability disclosure. Hence, Prometheus performs similarity matching among vulnerabilities within the same layer and across different layers, revealing how vulnerabilities in one layer can lead to the exploitation of vulnerabilities in different layers. This approach allows for prioritized risk analysis, mitigation strategies, and patching efforts based on the specific nature and severity of vulnerabilities at each layer.
@@
As a result, Prometheus automatically produces two types of attack graphs: cumulative (or multi-layer) attack graphs and layered attack graphs. Cumulative attack graphs show how an attacker could exploit Common Vulnerabilities and Exposures (CVEs) across multiple layers. In contrast, layered attack graphs focus on the exploitation of vulnerabilities within the same layer. This dual representation provides a comprehensive understanding of the potential attack paths across different layers and within individual layers.
@@
By traversing the generated graphs, Prometheus thoroughly analyzes the security posture of the entire infrastructure, providing both cumulative and layer-wise security analytics. These analytics reflect (1) the exploitability efforts required by an adversary for a step or a sequence of steps and (2) the impact of attack steps based on the criticality of the resources affected in the analyzed infrastructure. For instance, resources considered critical in a specific deployment substantially increase the impact of an attack. Once scores are computed with respect to such criticality, Prometheus identifies vulnerabilities or highly exploitable and high-impact attack paths that require immediate attention to safeguard the system from potential attacks.
@@
Section 2 motivates Prometheus. Section 3 provides background information on vulnerability disclosures and on natural language processing using NER. Section 4 gives an overview of the Prometheus pipeline, while in Section 5, we delve into the technical details, starting from the ML-based processing to the attack graph generation. Subsequently, Section 6 outlines the risk analysis methods incorporated in Prometheus, designed to evaluate the security posture of a given infrastructure. In Section 7, we provide a comprehensive discussion of the implementation details, followed by the presentation of the evaluation results in Section 8. Finally, Section 9 discusses the related work and Section 10 concludes the paper.
@@ 2. Motivation
In this section, we delve into the rationale behind employing attack graphs as a fundamental tool for conducting posture analysis. We then discuss limitations of previous approaches - limitations that are addressed by Prometheus.
@@ 4. System Overview
In this section, we present an overview of the Prometheus pipeline, followed by a discussion of the system’s architecture.
Figure 2 provides a comprehensive overview of the Prometheus pipeline, encompassing the following key steps:
@@
(3) Attack Graph Construction: In this phase, Prometheus constructs attack graphs by integrating the attack graph nodes identified in step 1 with the vulnerability similarity results obtained in step 2.
(4) Risk Analysis: Finally, Prometheus conducts a thorough security posture analysis based on the attack graphs generated by the previous step. Prometheus traverses the graphs, utilizing various vulnerability parameters such as exploitability and risk scores. This quantification process is essential for assessing the security posture layer by layer and cumulatively. Additionally, it helps identify the most impactful attack paths within the network infrastructure.
Figure 2. Prometheus’ pipeline.
@@ 5.2. Attack Graph Edge Connection
The key task for the word embeddings for Prometheus is to curate the security corpus that is general to different vulnerabilities and attack scenarios. As discussed in Section 3, the NVD database contains detailed vulnerability descriptions and references of CVEs that have undergone thorough manual assessment and processing (booth2013national, ). This has resulted in the generation of succinct information regarding various vulnerabilities. For the CVE description collection, we download CVE feeds hosted on the NVD website, which package CVE entries with descriptions. Moreover, there are also many other online resources, such as MITRE ATT&CK knowledge base (strom2018mitre, ), that collect comprehensive vulnerability information such as Common Vulnerability and Exposures (CWEs), which we also consider in our dataset curation process.
@@ 5.3. Attack Graph Construction and Partition
The attack graph comprises three primary node types: the attacker (source nodes), existing CVEs associated with different network entities (intermediate nodes), and CWEs serving as attacking targets (sink nodes). The construction of edges connecting nodes and the assignment of corresponding edge weights depend on the types of connected nodes and available data. By default, an edge is established from the attacker node to each CVE node, assuming the attacker can exploit the CVE. This inclusivity aligns with Prometheus’ objective to generate all conceivable scenarios under various adversarial assumptions in the network. For example, if a CVE’s attack vector necessitates physical access to the device, Prometheus includes an analysis of such a scenario. Users can subsequently filter the generated attack graphs based on their network’s specific adversarial assumptions. The edge weights for these edges are predominantly determined by CVSS base scores, indicating the likelihood of a successful attacker exploit, directly correlating with the severity of the threat posed by exposure to such vulnerabilities.
@@ 91 XML Injection (aka Blind XPath Injection)
Prometheus follows two approaches to classify the identified vulnerabilities to the layers of interest. In the first approach, Prometheus uses a predefined set of high-frequency keywords that typically appear in the corresponding layer. Table 1 shows a snippet of the keywords Prometheus uses for mapping the vulnerabilities to the network layer. We list the complete list of keywords in Appendix A. The second approach is based on the MITRE Common Weakness Enumeration (CWE) identifier appointed by the vulnerability database. Table 2 shows a snippet of the list of CWEs Prometheus uses for mapping the vulnerabilities to the network layer. We list the complete list in Appendix A. The reason for using both approaches is that, on the one hand, only keyword matching may result in misclassified vulnerabilities. On the other hand, relying merely on the CWEs is not enough, as many vulnerabilities have not been appointed to CWEs. Additionally, the Dashboard service allows an administrator to change the layer to which a particular vulnerability has been assigned.
@@ 6. Risk Scoring System
Within the graphs, each node (i.e., CVE) is linked to exploitability, impact, and risk scores. These scores are derived in accordance with the Common Vulnerability Scoring System (CVSS) standard (cvss3x, ) (version 3.1). Recognizing that these scores offer a limited, isolated perspective on vulnerability impact, Prometheus enhances the computed scores based on CVSS standards and integrates them for the risk assessment of attack graphs, considering the criticality of affected resources. The evaluation of security posture involves:
@@
The functions $eScore()$ and $iScore()$ within Prometheus are intentionally designed to be user-provided. Formulating a precise function that accurately encapsulates the exploitability and impact scores across diverse deployment scenarios poses a considerable challenge. For instance, the impact of a CVE affecting a device deployed within a critical infrastructure network significantly differs from its impact when the same device is located in a publicly-facing demilitarized zone. To address this complexity, users are granted the flexibility to define their own customized functions. These functions can incorporate the assigned CVE scores along with other parameters of their choosing, allowing for a more tailored and context-aware evaluation of exploitability and impact in varied deployment scenarios. If those functions are not provided, Prometheus uses each CVE’s assigned CVSS exploitability and impact scores.
@@
Graph Scores. For each generated graph, Prometheus computes the graph exploit, impact, and risk scores. Each score is computed as the average of the EES, EIS, and ERS sets.
@@
Identifying the high severity attack paths. To compute all the paths from the attacker node to the attacker’s goals, Prometheus searches for all the possible paths up to a certain number of edges; this is the cutoff limit for path explorations. For each computed path, Prometheus computes the exploitability, impact, and risk score of every path, which is essentially the total sum of EES, EIS, and ERS of the edges in the path. Then Prometheus sorts the paths in descending order of risk, exploitability, and impact. Prometheus allows the system administrator to change how sorting is performed through the dashboard interface.
@@
Identifying the minimum set of vulnerabilities that cover all the attack paths. We apply a minimum set of vertex cover on the constructed attack graph to identify the minimum set of vulnerabilities that cover all the edges in the graph. As vertex cover is an NP-hard problem, we use a local-ratio approximation algorithm to find the minimum set vertex cover (bar1985local, ). Thus, Prometheus identifies the minimum set of vulnerabilities that could render the attack paths impractical once mitigated.
@@ 7. Implementation
Figure 6. Prometheus is composed of five microservices (MS): the Dashboard MS, Graph MS, Machine Learning MS, Risk Scoring MS, and Database MS.
We have implemented Prometheus of 6,116 lines of Python codes. Prometheus is comprised of five microservices (MS) (see  Figure 6). Steps 1 and 2 of the pipeline (see Figure 2) are implemented by the Machine Learning MS, step 3 by Graph MS and step 4 by the Risk Scoring MS. Those services are implemented with 987 lines of Python code. Specifically, we build the named entity recognition and word embedding models based on spaCy (honnibal2020spacy, ), Gensim (rehurek2011gensim, ), and Scikit-learn (pedregosa2011scikit, ). For the risk scoring system, we use the NetworkX (hagberg2008exploring, ) library for graph construction, graph traversal, and risk score computation.
The Dashboard MS orchestrates the entire Prometheus pipeline. It has been developed with about 1.5K lines of code using Flask (flask, ) and Dash (dash, ) libraries. The dashboard provides an interface allowing the user (e.g., security officers and network admins) to import details about the network infrastructure, such as the network topology, communicating entities, and device details. Once the posture analysis is completed, the dashboard presents the generated graphs and the results of the risk analysis. For instance, Figure 8 shows Prometheus’ front-end dashboard interface snapshot. The figure showcases a spider graph delineating the top three identified vulnerabilities and subsequently outlines the top three high-severity paths. Additionally, a table excerpt is presented, detailing identified vulnerabilities along with their corresponding extracted pre- and postconditions.
@@
The user can also introduce new nodes in the graph (i.e., new CVEs) and modify or delete existing ones. Such functionality is needed as often online resources are not immediately updated once a certain vulnerability is disclosed. Moreover, the risk, exploitability, and impact scores may not reflect the criticality of the resources in the infrastructure (e.g., overestimated or underestimated scores). Thus, Prometheus allows users to change those scores. Once the user makes such changes, the dashboard communicates the request to the back end, where a change-impact analysis occurs. The results are returned to the dashboard, which is immediately updated.
Figure 8. A small snapshot of Prometheus’ front-end dashboard interface.
@@
The dashboard also provides visual representations of the generated attack graphs pertaining to distinct layers within the designated infrastructure. For instance, Figure 7 shows a multi-layer attack graph generated by Prometheus. In the example in the figure, the initial CVE node is categorized within the Machine Learning (ML) layer, while the subsequent CVE node is categorized within the domain of hardware and systems.

### v2

# Graphene: Infrastructure Security Posture Analysis with AI-generated Attack Graphs

@@ Abstract.
The rampant occurrence of cybersecurity breaches imposes substantial limitations on the progress of network infrastructures, leading to compromised data, financial losses, potential harm to individuals, and disruptions in essential services. The current security landscape demands the urgent development of a holistic security assessment solution that encompasses vulnerability analysis and investigates the potential exploitation of these vulnerabilities as attack paths. In this paper, we propose Graphene, an advanced system designed to provide a detailed analysis of the security posture of computing infrastructures. Using user-provided information, such as device details and software versions, Graphene performs a comprehensive security assessment. This assessment includes identifying associated vulnerabilities and constructing potential attack graphs that adversaries can exploit. Furthermore, Graphene evaluates the exploitability of these attack paths and quantifies the overall security posture through a scoring mechanism. The system takes a holistic approach by analyzing security layers encompassing hardware, system, network, and cryptography. Furthermore, Graphene delves into the interconnections between these layers, exploring how vulnerabilities in one layer can be leveraged to exploit vulnerabilities in others. In this paper, we present the end-to-end pipeline implemented in Graphene, showcasing the systematic approach adopted for conducting this thorough security analysis.
@@ 1. Introduction
Our Approach. To address such challenges, we design an innovative solution, called Graphene, which serves as a comprehensive security posture analyzer for computing infrastructures and applications. Graphene operates seamlessly by continuously monitoring trustworthy data sources, such as national vulnerability databases, for disclosed vulnerabilities specific to the devices and configuration of a given infrastructure.
@@
Using named entity recognition (NER), Graphene is able to extract the semantic meaning of these vulnerabilities and encode them into a latent space for in-depth analysis. In particular, Graphene automatically extracts the preconditions required for an adversary to exploit a vulnerability and the result after exploiting the vulnerability (i.e., postconditions). For example, Graphene can discern that exploiting a particular CVE might require the presence of the TensorFlow XLA compiler in the Google TensorFlow version prior to 1.7.0 as a precondition. Simultaneously, Graphene captures postconditions such as a system crash resulting from a heap buffer overflow or reading from other parts of the process memory.
@@
Subsequently, Graphene employs a semantic similarity-matching approach based on word embeddings (li2018word, ) to evaluate whether the postconditions of a vulnerability match the preconditions of another, thus constructing potential attack graphs for a given topology. Once generated, Graphene analyzes these attack graphs to evaluate the security posture. We propose a novel layered approach to this analysis, where vulnerabilities within each layer share similar characteristics and can be comprehensively assessed by subject matter experts. Graphene categorizes vulnerabilities in the generated attack graphs into distinct layers, such as machine learning, system, hardware, network, and cryptography, using keyword matching and common weakness enumeration (CWE) information within the vulnerability disclosure. Hence, Graphene performs similarity matching among vulnerabilities within the same layer and across different layers, revealing how vulnerabilities in one layer can lead to the exploitation of vulnerabilities in different layers. This approach allows for prioritized risk analysis, mitigation strategies, and patching efforts based on the specific nature and severity of vulnerabilities at each layer.
@@
As a result, Graphene automatically produces two types of attack graphs: cumulative (or multi-layer) attack graphs and layered attack graphs. Cumulative attack graphs show how an attacker could exploit Common Vulnerabilities and Exposures (CVEs) across multiple layers. In contrast, layered attack graphs focus on the exploitation of vulnerabilities within the same layer. This dual representation provides a comprehensive understanding of the potential attack paths across different layers and within individual layers.
@@
By traversing the generated graphs, Graphene thoroughly analyzes the security posture of the entire infrastructure, providing both cumulative and layer-wise security analytics. These analytics reflect (1) the exploitability efforts required by an adversary for a step or a sequence of steps and (2) the impact of attack steps based on the criticality of the resources affected in the analyzed infrastructure. For instance, resources considered critical in a specific deployment substantially increase the impact of an attack. Once scores are computed with respect to such criticality, Graphene identifies vulnerabilities or highly exploitable and high-impact attack paths that require immediate attention to safeguard the system from potential attacks.
@@
Section 2 motivates Graphene. Section 3 provides background information on vulnerability disclosures and on natural language processing using NER. Section 4 gives an overview of the Graphene pipeline, while in Section 5, we delve into the technical details, starting from the ML-based processing to the attack graph generation. Subsequently, Section 6 outlines the risk analysis methods incorporated in Graphene, designed to evaluate the security posture of a given infrastructure. In Section 7, we provide a comprehensive discussion of the implementation details, followed by the presentation of the evaluation results in Section 8. Finally, Section 9 discusses the related work and Section 10 concludes the paper.
@@ 2. Motivation
In this section, we delve into the rationale behind employing attack graphs as a fundamental tool for conducting posture analysis. We then discuss limitations of previous approaches - limitations that are addressed by Graphene.
@@ 4. System Overview
In this section, we present an overview of the Graphene pipeline, followed by a discussion of the system’s architecture.
Figure 2 provides a comprehensive overview of the Graphene pipeline, encompassing the following key steps:
@@
(3) Attack Graph Construction: In this phase, Graphene constructs attack graphs by integrating the attack graph nodes identified in step 1 with the vulnerability similarity results obtained in step 2.
(4) Risk Analysis: Finally, Graphene conducts a thorough security posture analysis based on the attack graphs generated by the previous step. Graphene traverses the graphs, utilizing various vulnerability parameters such as exploitability and risk scores. This quantification process is essential for assessing the security posture layer by layer and cumulatively. Additionally, it helps identify the most impactful attack paths within the network infrastructure.
Figure 2. Graphene’ pipeline.
@@ 5.2. Attack Graph Edge Connection
The key task for the word embeddings for Graphene is to curate the security corpus that is general to different vulnerabilities and attack scenarios. As discussed in Section 3, the NVD database contains detailed vulnerability descriptions and references of CVEs that have undergone thorough manual assessment and processing (booth2013national, ). This has resulted in the generation of succinct information regarding various vulnerabilities. For the CVE description collection, we download CVE feeds hosted on the NVD website, which package CVE entries with descriptions. Moreover, there are also many other online resources, such as MITRE ATT&CK knowledge base (strom2018mitre, ), that collect comprehensive vulnerability information such as Common Vulnerability and Exposures (CWEs), which we also consider in our dataset curation process.
@@ 5.3. Attack Graph Construction and Partition
The attack graph comprises three primary node types: the attacker (source nodes), existing CVEs associated with different network entities (intermediate nodes), and CWEs serving as attacking targets (sink nodes). The construction of edges connecting nodes and the assignment of corresponding edge weights depend on the types of connected nodes and available data. By default, an edge is established from the attacker node to each CVE node, assuming the attacker can exploit the CVE. This inclusivity aligns with Graphene’ objective to generate all conceivable scenarios under various adversarial assumptions in the network. For example, if a CVE’s attack vector necessitates physical access to the device, Graphene includes an analysis of such a scenario. Users can subsequently filter the generated attack graphs based on their network’s specific adversarial assumptions. The edge weights for these edges are predominantly determined by CVSS base scores, indicating the likelihood of a successful attacker exploit, directly correlating with the severity of the threat posed by exposure to such vulnerabilities.
@@ 91 XML Injection (aka Blind XPath Injection)
Graphene follows two approaches to classify the identified vulnerabilities to the layers of interest. In the first approach, Graphene uses a predefined set of high-frequency keywords that typically appear in the corresponding layer. Table 1 shows a snippet of the keywords Graphene uses for mapping the vulnerabilities to the network layer. We list the complete list of keywords in Appendix A. The second approach is based on the MITRE Common Weakness Enumeration (CWE) identifier appointed by the vulnerability database. Table 2 shows a snippet of the list of CWEs Graphene uses for mapping the vulnerabilities to the network layer. We list the complete list in Appendix A. The reason for using both approaches is that, on the one hand, only keyword matching may result in misclassified vulnerabilities. On the other hand, relying merely on the CWEs is not enough, as many vulnerabilities have not been appointed to CWEs. Additionally, the Dashboard service allows an administrator to change the layer to which a particular vulnerability has been assigned.
@@ 6. Risk Scoring System
Within the graphs, each node (i.e., CVE) is linked to exploitability, impact, and risk scores. These scores are derived in accordance with the Common Vulnerability Scoring System (CVSS) standard (cvss3x, ) (version 3.1). Recognizing that these scores offer a limited, isolated perspective on vulnerability impact, Graphene enhances the computed scores based on CVSS standards and integrates them for the risk assessment of attack graphs, considering the criticality of affected resources. The evaluation of security posture involves:
@@
The functions $eScore()$ and $iScore()$ within Graphene are intentionally designed to be user-provided. Formulating a precise function that accurately encapsulates the exploitability and impact scores across diverse deployment scenarios poses a considerable challenge. For instance, the impact of a CVE affecting a device deployed within a critical infrastructure network significantly differs from its impact when the same device is located in a publicly-facing demilitarized zone. To address this complexity, users are granted the flexibility to define their own customized functions. These functions can incorporate the assigned CVE scores along with other parameters of their choosing, allowing for a more tailored and context-aware evaluation of exploitability and impact in varied deployment scenarios. If those functions are not provided, Graphene uses each CVE’s assigned CVSS exploitability and impact scores.
@@
Graph Scores. For each generated graph, Graphene computes the graph exploit, impact, and risk scores. Each score is computed as the average of the EES, EIS, and ERS sets.
@@
Identifying the high severity attack paths. To compute all the paths from the attacker node to the attacker’s goals, Graphene searches for all the possible paths up to a certain number of edges; this is the cutoff limit for path explorations. For each computed path, Graphene computes the exploitability, impact, and risk score of every path, which is essentially the total sum of EES, EIS, and ERS of the edges in the path. Then Graphene sorts the paths in descending order of risk, exploitability, and impact. Graphene allows the system administrator to change how sorting is performed through the dashboard interface.
@@
Identifying the minimum set of vulnerabilities that cover all the attack paths. We apply a minimum set of vertex cover on the constructed attack graph to identify the minimum set of vulnerabilities that cover all the edges in the graph. As vertex cover is an NP-hard problem, we use a local-ratio approximation algorithm to find the minimum set vertex cover (bar1985local, ). Thus, Graphene identifies the minimum set of vulnerabilities that could render the attack paths impractical once mitigated.
@@ 7. Implementation
Figure 6. Graphene is composed of five microservices (MS): the Dashboard MS, Graph MS, Machine Learning MS, Risk Scoring MS, and Database MS.
We have implemented Graphene of 6,116 lines of Python codes. Graphene is comprised of five microservices (MS) (see  Figure 6). Steps 1 and 2 of the pipeline (see Figure 2) are implemented by the Machine Learning MS, step 3 by Graph MS and step 4 by the Risk Scoring MS. Those services are implemented with 987 lines of Python code. Specifically, we build the named entity recognition and word embedding models based on spaCy (honnibal2020spacy, ), Gensim (rehurek2011gensim, ), and Scikit-learn (pedregosa2011scikit, ). For the risk scoring system, we use the NetworkX (hagberg2008exploring, ) library for graph construction, graph traversal, and risk score computation.
The Dashboard MS orchestrates the entire Graphene pipeline. It has been developed with about 1.5K lines of code using Flask (flask, ) and Dash (dash, ) libraries. The dashboard provides an interface allowing the user (e.g., security officers and network admins) to import details about the network infrastructure, such as the network topology, communicating entities, and device details. Once the posture analysis is completed, the dashboard presents the generated graphs and the results of the risk analysis. For instance, Figure 8 shows Graphene’ front-end dashboard interface snapshot. The figure showcases a spider graph delineating the top three identified vulnerabilities and subsequently outlines the top three high-severity paths. Additionally, a table excerpt is presented, detailing identified vulnerabilities along with their corresponding extracted pre- and postconditions.
@@
The user can also introduce new nodes in the graph (i.e., new CVEs) and modify or delete existing ones. Such functionality is needed as often online resources are not immediately updated once a certain vulnerability is disclosed. Moreover, the risk, exploitability, and impact scores may not reflect the criticality of the resources in the infrastructure (e.g., overestimated or underestimated scores). Thus, Graphene allows users to change those scores. Once the user makes such changes, the dashboard communicates the request to the back end, where a change-impact analysis occurs. The results are returned to the dashboard, which is immediately updated.
Figure 8. A small snapshot of Graphene’ front-end dashboard interface.
@@
The dashboard also provides visual representations of the generated attack graphs pertaining to distinct layers within the designated infrastructure. For instance, Figure 7 shows a multi-layer attack graph generated by Graphene. In the example in the figure, the initial CVE node is categorized within the Machine Learning (ML) layer, while the subsequent CVE node is categorized within the domain of hardware and systems.